OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910 I NEED HELP I Have 2 More Problems To Answer PLEASE HELP

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = −1. Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y=-1 Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Squirrels @dan815

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

@Cacciatore_J

OpenStudy (anonymous):

@Donblue22 Are You Working On It?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

OMG This Website Is A JOKE!

OpenStudy (anonymous):

I'm here

OpenStudy (anonymous):

Sorry, was busy irl

OpenStudy (anonymous):

Thank God! Lol

OpenStudy (anonymous):

One moment, I'm refreshing my memory *hasn't done this in a while* I'll be right with you

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So, your focus is 1, where y=term a in the general form of a parabola: y^2=4ax or x^2=4ay

OpenStudy (anonymous):

The Answer Should Be Something Like f(x)=x over x x^2 X Will Be Replace By Numbers Of

OpenStudy (anonymous):

Course

OpenStudy (anonymous):

There is an equation button just below the comment box. I want to be sure that I interpreted that result of yours properly, so would you mind typing it into that?

OpenStudy (anonymous):

Like That

OpenStudy (anonymous):

Why Wont It Let Me Insert

OpenStudy (anonymous):

\[f(x)=x/(x(x^2))\]

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

Its f(x)= (fraction) (x- or + 2)^2+6 or 8

OpenStudy (anonymous):

From who or where did you get this equation(s)? What makes you think that this is the answer?

OpenStudy (anonymous):

FLVS

OpenStudy (anonymous):

Okay, so they'll want the answer in that format... The thing is, that doesn't resemble the equation that I believe it to be. And I don't know if I can cater my formula into that format, simply because the numbers don't look right. Would you like to try my method, or find someone from FLVS to do it in your method?

OpenStudy (anonymous):

Try Yours

OpenStudy (anonymous):

Alright then, I'll keep it as simple as possible. The vertex isn't given,but since there aren't any intercepts or such given, it's safe to assume that the vertex is the origin Because the focus is 1, and the directrix is -1, we know that p from x^2=4py is equal to 1

OpenStudy (anonymous):

The directrix is a line never touched by the line of the equation, so we know that the equation opens upwards